A 20 kg traveler's bag is being pulled along the floor by tourist with a force of 50N. The force is applied on the handle that forms an angle 30 ^ 0 with horizontal . Neglecting friction, what is the acceleration of the bag? How much force is exerted by the floor on the bag, if friction is neglected
1
Expert's answer
2020-12-10T11:10:43-0500
"ma=F\cos{30}\\20a=50\cos{30}\\a=2.165\frac{m}{s^2}"

"N=mg-F\sin{30}\\=(20)(9.8)-50\sin{30}=171\ N"